# Distributional Reduction: Unifying Dimensionality Reduction and Clustering with Gromov-Wasserstein

### v2

@@ 1 Introduction
One major objective of unsupervised learning Hastie2009 is to provide interpretable and meaningful approximate representations of the data that best preserve its structure i.e. the underlying geometric relationships between the data samples. Similar in essence to Occam’s principle frequently employed in supervised learning, the preference for unsupervised data representation often aligns with the pursuit of simplicity, interpretability or visualizability in the associated model. These aspects are determinant in many real-world applications where the interaction with domain experts is paramount for interpreting the results and extracting meaningful insights from the model. For instance, the design of tissue atlases or the inference of single-cell trajectories, which are essential for addressing various biological challenges rao2021exploring ; saelens2019comparison , depends on the analysis of metacells, i.e. granular and interpretable representations of cells that enable separating sampling effect from biological variance baran2019metacell .
Dimensionality reduction and clustering. When faced with the question of extracting interpretable representations, from a dataset ${\mathbf{X}}=({\mathbf{x}}_{1},...,{\mathbf{x}}_{N})^{\top}\in\mathbb{R}^{N\times p}$ of $N$ samples in $\mathbb{R}^{p}$, the machine learning community has proposed a variety of methods. Among them, dimensionality reduction (DR) algorithms have been widely used to summarize data in a low-dimensional space ${\mathbf{Z}}=({\mathbf{z}}_{1},...,{\mathbf{z}}_{N})^{\top}\in\mathbb{R}^{N\times d}$ with $d\ll p$, allowing for visualization of every individual points for small enough $d$ lee2007nonlinear ; van2009dimensionality . Another major approach is to cluster the data into $n$ groups, with $n$ typically much smaller than $N$, and to summarize these groups through their centroids bishop2006pattern ; von2007tutorial . Clustering is particularly interpretable since it provides a smaller number of representative points that can be easily inspected. The cluster assignments can also be analyzed. Both DR and clustering follow a similar philosophy of summarization and reduction of the dataset using a smaller size representation.
Two sides of the same coin. As a matter of fact, methods from both families share many similitudes, including the construction of a similarity graph between input samples. In clustering, many popular approaches design a reduced or coarsened version of the initial similarity graph while preserving some of its spectral properties von2007tutorial ; schaeffer2007graph . In DR, the goal is to solve the inverse problem of finding low-dimensional embeddings that generate a similarity graph close to the one computed from input data points ham2004kernel ; hinton2002stochastic . Our work builds on these converging viewpoints and addresses the following question: can DR and clustering be expressed in a common and unified framework ?
@@
Figure 1: Illustration of our DistR method on two toys examples. Both pairs of representations ${\mathbf{X}},{\mathbf{Z}}$ are respectively associated to their similarity matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$, coupled via the matrix ${\mathbf{T}}$ (depicted in purple, with its marginals).
@@
A distributional perspective. To answer this question, we propose to look at both problems from a distributional point of view, treating the data as an empirical probability distribution $\mu=\frac{1}{N}\sum_{i}\delta_{{\mathbf{x}}_{i}}$. This enables to consider statistical measures of similarity such as Optimal Transport (OT), which is at the core of our work. On the one hand, OT and clustering are strongly related. The celebrated k-means algorithm can be seen as a particular case of minimal Wasserstein estimator where a distribution of $n$ Diracs is optimized w.r.t their weights and positions Canas12 . Other connections between spectral clustering and the OT-based Gromov-Wasserstein (GW) distance have been recently developed in chowdhury2021generalized ; chen2023gromov ; vincent2021semi . On the other hand, the link between DR and OT has not been explored yet. DR methods, when modeling data as distributions, usually focus on joint distribution between samples within each space separately, see e.g. van2023snekhorn or lu2019doubly . Consequently, they do not consider couplings to transport samples across spaces of varying dimensions.
@@
Contributions. In this paper, we bridge this gap by proposing a general distributional framework that encompasses both DR and clustering as special cases. We notably cast those problems as finding a reduced distribution that minimizes the GW divergence from the original empirical data distribution. Our method proceeds by first constructing an input similarity matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ that is matched with the embedding similarity ${\mathbf{C}}_{Z}({\mathbf{Z}})$ through an OT coupling matrix ${\mathbf{T}}$. The latter establishes correspondences between input and embedding samples. We illustrate this principle in Figure 1 where one can notice that ${\mathbf{C}}_{Z}({\mathbf{Z}})$ preserves the topology of ${\mathbf{C}}_{X}({\mathbf{X}})$ with a reduced number of nodes. The adaptivity of our model that can select an effective number of cluster $<n$, is visible in the bottom plot, where only the exact number of clusters in the original data ($9$ out of the $12$ initially proposed) is automatically recovered. Our method can operate in any embedding space, which is illustrated by projecting in either a 2D Euclidean plane or a Poincaré ball as embedding spaces.
@@
Outline. We show that this framework is versatile and allows recovering many popular DR methods such as the kernel PCA and neighbor embedding algorithms, but also clustering algorithms such as weighted k-means and its kernel counterpart including spectral clustering chan2004optimization ; dhillon2007weighted . We first prove in Section 3 that DR can be formulated as a GW projection problem under some conditions on the loss and similarity functions. We then propose in Section 4 a novel formulation of data summarization as a minimal GW estimator that allows selecting both the dimensionality of the embedding $d$ (DR) and the cardinality of the support $n$ (Clustering). Finally, we show in section 5 the practical interest of our approach on images and genomics datasets.
@@ 2.1 Unified View of Dimensionality Reduction
where $\|\cdot\|_{F}$ is the Frobenius norm. This problem is commonly known as kernel Principal Component Analysis (PCA) scholkopf1997kernel and an optimal solution is given by ${\mathbf{Z}}^{\star}=(\sqrt{\lambda_{1}}{\mathbf{v}}_{1},...,\sqrt{\lambda_{d}}{\mathbf{v}}_{d})^{\top}$ where $\lambda_{i}$ is the $i$-th largest eigenvalue of ${\mathbf{C}}_{X}({\mathbf{X}})$ with corresponding eigenvector ${\mathbf{v}}_{i}$ eckart1936approximation . As shown by ham2004kernel ; ghojogh2021unified , numerous dimension reduction methods can be categorized in this manner. This includes PCA when ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{X}}{\mathbf{X}}^{\top}$ is the matrix of inner products in the input space; (classical) multidimensional scaling borg2005modern , when ${\mathbf{C}}_{X}({\mathbf{X}})=-\frac{1}{2}{\mathbf{H}}{\mathbf{D}}_{\mathbf{X}}{\mathbf{H}}$ with ${\mathbf{D}}_{\mathbf{X}}$ the matrix of squared euclidean distance between the points in $\mathbb{R}^{p}$ and ${\mathbf{H}}={\mathbf{I}}_{N}-\frac{1}{N}\bm{1}_{N}\bm{1}_{N}^{\top}$ is the centering matrix; Isomap tenenbaum2000global , with ${\mathbf{C}}_{X}({\mathbf{X}})=-\frac{1}{2}{\mathbf{H}}{\mathbf{D}}_{\mathbf{X}}^{(g)}{\mathbf{H}}$ with ${\mathbf{D}}_{\mathbf{X}}^{(g)}$ the geodesic distance matrix; Laplacian Eigenmap belkin2003laplacian , with ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{L}}_{\mathbf{X}}^{\dagger}$ the pseudo-inverse of the Laplacian associated to some adjacency matrix ${\mathbf{W}}_{\mathbf{X}}$; but also Locally Linear Embedding roweis2000nonlinear , and Diffusion Map coifman2006diffusion (for all of these examples we refer to (ghojogh2021unified, , Table 1)).
@@
Within our framework, this corresponds to eq. DR with $L=L_{\operatorname{KL}}$. The objective function of popular methods such as stochastic neighbor embedding (SNE) hinton2002stochastic or t-SNE van2008visualizing can be derived from eq. NE with a particular choice of ${\mathbf{C}}_{X},{\mathbf{C}}_{Z}$. For instance SNE and t-SNE both consider in the input space a symmetrized version of the entropic affinity vladymyrov2013entropic ; van2023snekhorn . In the embedding space, ${\mathbf{C}}_{Z}({\mathbf{Z}})$ is usually constructed from a “kernel” matrix ${\mathbf{K}}_{\mathbf{Z}}$ which undergoes a scalar van2008visualizing , row-stochastic hinton2002stochastic or doubly stochastic lu2019doubly ; van2023snekhorn normalization. Gaussian kernel $[{\mathbf{K}}_{\mathbf{Z}}]_{ij}=\exp(-\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})$, or heavy-tailed Student-t kernel $[{\mathbf{K}}_{\mathbf{Z}}]_{ij}=(1+\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})^{-1}$, are typical choices van2008visualizing . We also emphasize that one can retrieve the UMAP objective mcinnes2018umap from eq. DR using the binary cross-entropy loss. A comprehensive overview and probabilistic analysis of these methods can be found in van2022probabilistic .
@@
Non-Euclidean geometries. Most DR methods can also be extended to incorporate non-Euclidean geometries. For instance, Hyperbolic spaces Chami21 ; Fan_2022_CVPR ; Guo22 ; Lin23 are of particular interest as they can capture hierarchical structures more effectively than Euclidean spaces and further mitigate the curse of dimensionality. However, these methods introduce additional hyperparameters and optimization difficulties that can limit their applicability. For this reason, we report some numerical experiments with such kernels in Section E.7 as proofs of concept supporting the versatility of our methods.
@@ 2.2 Optimal Transport Across Spaces
Optimal Transport (OT) villani2009optimal ; peyre2019computational is a popular framework for comparing probability distributions and is at the core of our contributions. We review in this section the Gromov-Wasserstein formulation of OT aiming at comparing distributions across spaces.
Gromov-Wasserstein (GW). The GW framework memoli2011gromov ; sturm2012space comprises a collection of OT methods designed to compare distributions by examining the pairwise relations within each space. For two matrices ${\mathbf{C}}\in\mathbb{R}^{N\times N}$, $\overline{{\mathbf{C}}}\in\mathbb{R}^{n\times n}$, and weights ${\mathbf{h}}\in\Sigma_{N},\overline{{\mathbf{h}}}\in\Sigma_{n}$, the GW discrepancy with inner loss $L$ peyre2016gromov is defined as
$\begin{split}&\operatorname{GW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=\min_{{\mathbf{T}}\in{\mathcal{U}}({\mathbf{h}},\overline{{\mathbf{h}}})}E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})\ =\min_{{\mathbf{T}}\in{\mathcal{U}}({\mathbf{h}},\overline{{\mathbf{h}}})}\sum_{ijkl}L(C_{ij},\overline{C}_{kl})T_{ik}T_{jl}\,,\end{split}$ (GW)
@@
From a distributional perspective, GW can also be viewed as a distance between distributions that do not belong to the same metric space. For two discrete probability distributions $\mu_{X}=\sum_{i=1}^{N}[{\mathbf{h}}_{X}]_{i}\delta_{{\mathbf{x}}_{i}}\in{\mathcal{P}}_{N}(\mathbb{R}^{p}),\mu_{Z}=\sum_{i=1}^{n}[{\mathbf{h}}_{Z}]_{i}\delta_{{\mathbf{z}}_{i}}\in{\mathcal{P}}_{n}(\mathbb{R}^{d})$ and pairwise similarity matrices ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$ associated with the explicit supports ${\mathbf{X}}=({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{n})^{\top}$ and ${\mathbf{Z}}=({\mathbf{z}}_{1},\cdots,{\mathbf{z}}_{n})^{\top}$, the quantity $\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{h}}_{X},{\mathbf{h}}_{Z})$ is a measure of dissimilarity or discrepancy between $\mu_{X},\mu_{Z}$. Specifically, when $L=L_{2}$, and ${\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}})$ are pairwise distance matrices, GW defines a proper distance between $\mu_{X}$ and $\mu_{Z}$ with respect to measure preserving isometries [^1] With weaker assumptions on ${\mathbf{C}}_{X},{\mathbf{C}}_{Z}$, GW defines a pseudo-metric w.r.t. a different notion of isomorphism chowdhury2019gromov . See Appendix B for more details. memoli2011gromov .
Due to its versatile properties, notably in comparing distributions over different domains, the GW problem has found many applications in machine learning, e.g., for 3D meshes alignment solomon2016entropic ; ezuz2017gwcnn , NLP alvarez2018gromov , (co-)clustering peyre2016gromov ; redko2020co , single-cell analysis demetci2020gromov , neuroimaging thual2022aligning , graph representation learning xu2020gromov ; vincent2021online ; liu2022robust ; vincent2022template ; pmlr-v202-zeng23c and partitioning xu2019scalable ; chowdhury2021generalized . In this work, we leverage the GW discrepancy to extend classical DR approaches, framing them as the projection of a distribution onto a space of lower dimensionality.
3 Optimal Transport Formulation of DR

### v3

@@ 1 Introduction
One major objective of unsupervised learning (Hastie et al., 2009) is to provide interpretable and meaningful approximate representations of the data that best preserve its structure i.e. the underlying geometric relationships between the data samples. Similar in essence to Occam’s principle frequently employed in supervised learning, the preference for unsupervised data representation often aligns with the pursuit of simplicity, interpretability or visualizability in the associated model. These aspects are determinant in many real-world applications where the interaction with domain experts is paramount for interpreting the results and extracting meaningful insights from the model. For instance, the design of tissue atlases or the inference of single-cell trajectories, which are essential for addressing various biological challenges (Rao et al., 2021, Saelens et al., 2019), depends on the analysis of metacells, i.e. granular and interpretable representations of cells that enable separating sampling effect from biological variance (Baran et al., 2019).
Dimensionality reduction and clustering. When faced with the question of extracting interpretable representations, from a dataset ${\mathbf{X}}=({\mathbf{x}}_{1},...,{\mathbf{x}}_{N})^{\top}\in\mathbb{R}^{N\times p}$ of $N$ samples in $\mathbb{R}^{p}$, the machine learning community has proposed a variety of methods. Among them, dimensionality reduction (DR) algorithms have been widely used to summarize data in a low-dimensional space ${\mathbf{Z}}=({\mathbf{z}}_{1},...,{\mathbf{z}}_{N})^{\top}\in\mathbb{R}^{N\times d}$ with $d\ll p$, allowing for visualization of every individual points for small enough $d$ (Lee et al., 2007, Van Der Maaten et al., 2009). Another major approach is to cluster the data into $n$ groups, with $n$ typically much smaller than $N$, and to summarize these groups through their centroids (Bishop, 2006, Von Luxburg, 2007). Clustering is particularly interpretable since it provides a smaller number of representative points that can be more easily inspected. The cluster assignments can also be analyzed. Both DR and clustering follow a similar philosophy of summarization and reduction of the dataset using a smaller size representation.
Two sides of the same coin. As a matter of fact, methods from both families share many similitudes, including the construction of a similarity graph between input samples. In clustering, many popular approaches design a reduced or coarsened version of the initial similarity graph while preserving some of its spectral properties (Von Luxburg, 2007, Schaeffer, 2007). In DR, the goal is to solve the inverse problem of finding low-dimensional embeddings that generate a similarity graph close to the one computed from input data points (Ham et al., 2004, Hinton and Roweis, 2002). Our work builds on these converging viewpoints and addresses the following question: can DR and clustering be expressed in a common and unified framework ?
@@
Figure 1: Illustration of our DistR method on two toy examples. For each example (left and right), the data pairs ${\mathbf{X}}$ and ${\mathbf{Z}}$ represent the original data and their embedding as found by our method DistR. These pairs are associated to similarity matrices ${\mathbf{C}}_{X}({\mathbf{X}})\in\mathbb{R}^{N\times N}$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})\in\mathbb{R}^{n\times n}$ ($n<N$), which encode the pairwise similarities in their respective spaces. In the left example, we consider a Gaussian similarity for ${\mathbf{Z}}$ and Symmetric Entropic Affinity (SEA) (Van Assel et al., 2023) for ${\mathbf{X}}$. In the right example, we consider SEA for ${\mathbf{X}}$ and a similarity in the Lorentz model for ${\mathbf{Z}}$ (Nickel and Kiela, 2018). DistR also provides a coupling ${\mathbf{T}}$ between the points in ${\mathbf{X}}$ and ${\mathbf{Z}}$, illustrated in purple, with its marginals.
@@
A distributional perspective. To answer this question, we propose to look at both problems from a distributional point of view, treating the data as an empirical probability distribution $\mu=\frac{1}{N}\sum_{i}\delta_{{\mathbf{x}}_{i}}$. This enables to consider statistical measures of similarity such as Optimal Transport (OT), which is at the core of our work. On the one hand, OT and clustering are strongly related. The celebrated k-means algorithm can be seen as a particular case of minimal Wasserstein estimator where a distribution of $n$ Diracs is optimized w.r.t their weights and positions (Canas and Rosasco, 2012). Other connections between spectral clustering and the OT-based Gromov-Wasserstein (GW) distance have been recently developed in Chowdhury and Needham (2021), Chen et al. (2023), Vincent-Cuaz et al. (2022a). On the other hand, the link between DR and OT is less explored. DR methods, when modeling data as distributions, usually focus on joint distribution between samples within each space separately, see e.g. Van Assel et al. (2023) or Lu et al. (2019). Consequently, they do not consider couplings to transport samples across spaces of varying dimensions. At the time of this paper’s submission, we note that other authors have independently developed a similar line of work, as detailed in Clark et al. (2024), Murray and Pickarski (2024). Their studies focus on “multidimensional scaling” problems and establish connections to the semi-relaxed GW framework but do not explore alternative DR approaches or clustering aspects.
@@
Contributions. In this paper, we develop a general distributional framework that encompasses both DR and clustering as special cases. We notably cast those problems as finding a reduced distribution that minimizes the GW divergence from the original empirical data distribution. Our method proceeds by first constructing an input similarity matrix ${\mathbf{C}}_{X}({\mathbf{X}})$ that is matched with the embedding similarity ${\mathbf{C}}_{Z}({\mathbf{Z}})$ through an OT coupling matrix ${\mathbf{T}}$. The latter establishes correspondences between input and embedding samples. We illustrate this principle in Figure 1 where one can notice that ${\mathbf{C}}_{Z}({\mathbf{Z}})$ preserves the topology of ${\mathbf{C}}_{X}({\mathbf{X}})$ with a reduced number of nodes. The adaptivity of our model that can select an effective number of cluster $<n$, is visible in the bottom plot, where only the exact number of clusters in the original data ($9$ out of the $12$ initially proposed) is automatically recovered. Our method can operate in any embedding space, which is illustrated by projecting in either a 2D Euclidean plane or a Poincaré ball.
@@
Outline. We show that this framework is versatile and allows recovering as special cases many popular DR methods such as the kernel PCA and neighbor embedding algorithms, but also clustering algorithms such as weighted k-means and its kernel counterpart including spectral clustering (Chan et al., 2004, Dhillon et al., 2007). We first prove in Section 3 that DR can be formulated as a GW projection problem under some conditions on the loss and similarity functions. We then propose in Section 4 a novel formulation of data summarization as a minimal GW estimator that allows selecting both the dimensionality of the embedding $d$ (DR) and the cardinality of the support $n$ (Clustering). Finally, we show in section 5 the practical interest of our approach on images and genomics datasets.
@@ 2.1 Unified View of Dimensionality Reduction
where $\|\cdot\|_{F}$ is the Frobenius norm. This problem is commonly known as kernel Principal Component Analysis (PCA) (Schölkopf et al., 1997) and an optimal solution is given by ${\mathbf{Z}}^{\star}=(\sqrt{\lambda_{1}}{\mathbf{v}}_{1},...,\sqrt{\lambda_{d}}{\mathbf{v}}_{d})^{\top}$ where $\lambda_{i}$ is the $i$-th largest eigenvalue of ${\mathbf{C}}_{X}({\mathbf{X}})$ with corresponding eigenvector ${\mathbf{v}}_{i}$ (Eckart and Young, 1936). As shown by Ham et al. (2004), Ghojogh et al. (2021), numerous dimension reduction methods can be categorized in this manner. This includes PCA when ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{X}}{\mathbf{X}}^{\top}$ is the matrix of inner products in the input space; (classical) multidimensional scaling (Borg and Groenen, 2005), when ${\mathbf{C}}_{X}({\mathbf{X}})=-\frac{1}{2}{\mathbf{H}}{\mathbf{D}}_{\mathbf{X}}{\mathbf{H}}$ with ${\mathbf{D}}_{\mathbf{X}}$ the matrix of squared euclidean distance between the points in $\mathbb{R}^{p}$ and ${\mathbf{H}}={\mathbf{I}}_{N}-\frac{1}{N}\bm{1}_{N}\bm{1}_{N}^{\top}$ is the centering matrix; Isomap (Tenenbaum et al., 2000), with ${\mathbf{C}}_{X}({\mathbf{X}})=-\frac{1}{2}{\mathbf{H}}{\mathbf{D}}_{\mathbf{X}}^{(g)}{\mathbf{H}}$ with ${\mathbf{D}}_{\mathbf{X}}^{(g)}$ the geodesic distance matrix; Laplacian Eigenmap (Belkin and Niyogi, 2003), with ${\mathbf{C}}_{X}({\mathbf{X}})={\mathbf{L}}_{\mathbf{X}}^{\dagger}$ the pseudo-inverse of the Laplacian associated to some adjacency matrix ${\mathbf{W}}_{\mathbf{X}}$; but also Locally Linear Embedding (Roweis and Saul, 2000), and Diffusion Map (Coifman and Lafon, 2006) (for all of these examples we refer to Ghojogh et al. 2021, Table 1).
@@
Within our framework, this corresponds to eq. DR with $L=L_{\operatorname{KL}}$. The objective function of popular methods such as stochastic neighbor embedding (SNE) (Hinton and Roweis, 2002) or t-SNE (Van der Maaten and Hinton, 2008) can be derived from eq. NE with a particular choice of ${\mathbf{C}}_{X},{\mathbf{C}}_{Z}$. For instance SNE and t-SNE both consider in the input space a symmetrized version of the entropic affinity (Vladymyrov and Carreira-Perpinan, 2013, Van Assel et al., 2023). In the embedding space, ${\mathbf{C}}_{Z}({\mathbf{Z}})$ is usually constructed from a “kernel” matrix ${\mathbf{K}}_{\mathbf{Z}}$ which undergoes a scalar (Van der Maaten and Hinton, 2008), row-stochastic (Hinton and Roweis, 2002) or doubly stochastic (Lu et al., 2019, Van Assel et al., 2023) normalization. Gaussian kernel $[{\mathbf{K}}_{\mathbf{Z}}]_{ij}=\exp(-\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})$, or heavy-tailed Student-t kernel $[{\mathbf{K}}_{\mathbf{Z}}]_{ij}=(1+\|{\mathbf{z}}_{i}-{\mathbf{z}}_{j}\|_{2}^{2})^{-1}$, are typical choices (Van der Maaten and Hinton, 2008). We also emphasize that one can retrieve the UMAP objective (McInnes et al., 2018) from eq. DR using the binary cross-entropy loss $L_{BCE}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y}$. A comprehensive overview and probabilistic analysis of these methods can be found in Van Assel et al. (2022).
@@
Non-Euclidean geometries. Most DR methods can also be extended to incorporate non-Euclidean geometries. For instance, Hyperbolic spaces (Chami et al., 2021, Fan et al., 2022, Guo et al., 2022, Lin et al., 2023) are of particular interest as they can capture hierarchical structures more effectively than Euclidean spaces and further mitigate the curse of dimensionality. These methods introduce additional hyperparameters and optimization difficulties that can limit their applicability. For this reason, we report some numerical experiments with such kernels in Section F.10 as proofs of concept supporting the versatility of our method.
@@ 2.2 Optimal Transport Across Spaces
Optimal Transport (OT) (Villani et al., 2009, Peyré et al., 2019) is a popular framework for comparing probability distributions and is at the core of our contributions. We review in this section the Gromov-Wasserstein formulation of OT aiming at comparing distributions across spaces.
Gromov-Wasserstein (GW). The GW framework (Mémoli, 2011, Sturm, 2012) comprises a collection of OT methods designed to compare distributions by examining the pairwise relations within each space. For two matrices ${\mathbf{C}}\in\mathbb{R}^{N\times N}$, $\overline{{\mathbf{C}}}\in\mathbb{R}^{n\times n}$, and weights ${\mathbf{h}}\in\Sigma_{N},\overline{{\mathbf{h}}}\in\Sigma_{n}$, the GW discrepancy with inner loss $L$ (Peyré et al., 2016) is defined as
$\begin{split}&\operatorname{GW}_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{h}},\overline{{\mathbf{h}}})=\min_{{\mathbf{T}}\in{\mathcal{U}}({\mathbf{h}},\overline{{\mathbf{h}}})}E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})\ \\ &\text{with }E_{L}({\mathbf{C}},\overline{{\mathbf{C}}},{\mathbf{T}})=\sum_{ijkl}L(C_{ij},\overline{C}_{kl})T_{ik}T_{jl}\,,\end{split}\vspace{-2mm}$ (GW)
@@
From a distributional perspective, GW defines a distance between distributions that do not belong to the same metric space. For two discrete probability distributions $\mu_{X}=\sum_{i=1}^{N}[{\mathbf{h}}_{X}]_{i}\delta_{{\mathbf{x}}_{i}}\in{\mathcal{P}}_{N}(\mathbb{R}^{p}),\mu_{Z}=\sum_{i=1}^{n}[{\mathbf{h}}_{Z}]_{i}\delta_{{\mathbf{z}}_{i}}\in{\mathcal{P}}_{n}(\mathbb{R}^{d})$ and pairwise similarity matrices ${\mathbf{C}}_{X}({\mathbf{X}})$ and ${\mathbf{C}}_{Z}({\mathbf{Z}})$ associated with the explicit supports ${\mathbf{X}}=({\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{n})^{\top}$ and ${\mathbf{Z}}=({\mathbf{z}}_{1},\cdots,{\mathbf{z}}_{n})^{\top}$, the quantity $\operatorname{GW}_{L}({\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}}),{\mathbf{h}}_{X},{\mathbf{h}}_{Z})$ is a measure of dissimilarity or discrepancy between $\mu_{X},\mu_{Z}$. Specifically, when $L=L_{2}$, and ${\mathbf{C}}_{X}({\mathbf{X}}),{\mathbf{C}}_{Z}({\mathbf{Z}})$ are pairwise distance matrices, GW defines a proper distance between $\mu_{X}$ and $\mu_{Z}$ with respect to measure preserving isometries (with weaker assumptions on ${\mathbf{C}}_{X},{\mathbf{C}}_{Z}$, GW defines a pseudo-metric w.r.t. a different notion of isomorphism (Chowdhury and Mémoli, 2019), detailed in Appendix C).
Due to its versatile properties, notably in comparing distributions over different domains, the GW problem has found many applications in machine learning, e.g., for 3D meshes alignment (Solomon et al., 2016, Ezuz et al., 2017), NLP (Alvarez-Melis and Jaakkola, 2018), (co-)clustering (Peyré et al., 2016, Redko et al., 2020), single-cell analysis (Demetci et al., 2020), neuroimaging (Thual et al., 2022), graph representation learning (Xu, 2020, Vincent-Cuaz et al., 2021, Liu et al., 2022b, Vincent-Cuaz et al., 2022b, Zeng et al., 2023) and partitioning (Xu et al., 2019, Chowdhury and Needham, 2021). In this work, we leverage the GW discrepancy to extend classical DR approaches, framing them as the projection of a distribution onto a space of lower dimensionality.
3 OT Formulation of DR